How many ounces of pure salt must be added to 60 oz. of a 12% salt solution to make a 20% salt solutuion?
Answer by venugopalramana(3286)

How many ounces of pure salt must be added to 60 oz. of a 12% salt solution to make a 20% salt solutuion?
SOLUTION = 60 OZ.
ORIGINAL SALT CONTENT =12%....THAT IS 60*12/100=7.2 OZ
LET X OUNCES OF SALT BE ADDED TO GET THE REQD.SOLUTION
TOTAL SALT = 7.2+X OZ
SOLUTION = 60 OZ.
% COCN.= (7.2+X)100/60= 20%
7.2+X=20*60/100=12
X=12-7.2=4.8 OZ